OpenStudy (anonymous):

HELP PLEASE I GIVE MEDALS Multiply (x + 1) * (x -2)

OpenStudy (anonymous):

seems a bit easy I am in the 6th gradr i cant do higher

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

put the answers in so I can help you :)

OpenStudy (anonymous):

A. x^2 - 2x - 2 B. x^2 - x - 2 C. 2x - 2 D. x^2 -2

OpenStudy (anonymous):

alright cool. So is there ANY answer that you could possibly think it is?

OpenStudy (anonymous):

B or A seem possible

OpenStudy (anonymous):

Okay give me a couple seconds I'm gonna figure it out.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Alright so in math, for some stupid reason, we are always supposed to assume that x=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(1 + 1) * (1 - 2)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok so -1

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright so I hate this part probably more than you do :) you got your answer already ( it's -2) but they have all these weird answers that you have no idea what they mean. So go back to your original problem: (x + 1) * (x -2)

OpenStudy (anonymous):

ok what next?

OpenStudy (anonymous):

Alright so now distribute x*x. Then you do the x*2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand why you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I wish I could draw but my computer is going slow :/

OpenStudy (anonymous):

its ok no worries :)

OpenStudy (anonymous):

Lemme rephrase this whole thing... Here is your problem: (x + 1) * (x -2) Pull that first x over to the second x and multiply it. X*x= x^2 Then you pull the same x over to the 2 and you do x*2

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

So then you pull the 1 over to the x 1*x And that same 1 over to the 2 1*2 Do all of the steps and you'll have the answer :)

OpenStudy (anonymous):

omg is it A?

OpenStudy (anonymous):

No... :) Let me help you a little more.

OpenStudy (anonymous):

Step one: x*x= x^2 Step two: x*2 Step three: 1*x Step four: 1*2

OpenStudy (anonymous):

You already suggested the answer, and it's not A ;)

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

Yepperzz :) what grade are you in?

OpenStudy (anonymous):

9th :)

OpenStudy (anonymous):

Oh okay I'm in 8th hahaha

OpenStudy (anonymous):

thanks so much for your help this really helped :)

OpenStudy (anonymous):

No problem it helped me too haha :)

OpenStudy (anonymous):

BTW you're gorgeous!

OpenStudy (anonymous):

haha thanks but my profile picture is not me :) I just lost all my pictures on my laptop (hard drive crashed) but I look really similar to that so I just chose a picture off google

OpenStudy (anonymous):

ooh Okay haha I thought it was you I was that what the heeeelll

OpenStudy (anonymous):

my face + makeup is basic the same but my hair is blonde I'm considering dying it a brownish color next year :)

OpenStudy (anonymous):

and you're really pretty too

OpenStudy (anonymous):

aw thanks :)

OpenStudy (anonymous):

btw your bio says California school, like online school?

OpenStudy (anonymous):

yeah it's online but I have a teacher and classmates and i've met them all :) where do you live?

OpenStudy (anonymous):

New York I actually also go to an online school

OpenStudy (anonymous):

really? what's it called?

OpenStudy (anonymous):

Laurel Springs

OpenStudy (anonymous):

how cool is it K12 powered?

OpenStudy (anonymous):

yup :) i gtg finish some science assignments but it was great meeting youuu :) also if you have instagram or something tell me your user

OpenStudy (anonymous):

I don't have one :/ wayyyy to much drama but chat me on here? haha how lame is that :)

OpenStudy (anonymous):

haha its fine :)

OpenStudy (anonymous):

okay bye good luck with your schoolwork!

OpenStudy (anonymous):

ok ill message you later :)